Pleae help me solve this problem.


A right triangle has a height of x, a base of 5 and a hypotenuse of 13. Solve for x.

Answer by mananth(16946)

Use Pythagoras theorem
x^2 + 5^2= 13^2
x^2= 13^2-5^2
x^2=169-25
x^2=144
take the square root
x= 12 the height.
...
m.ananth@hotmail.ca